OpenStudy (anonymous):

Solve the equation 7 (d-5) + 12 = 5

OpenStudy (callisto):

7 (d-5) + 12 = 5 7d -35+12 =5 7d = 28 d=4

OpenStudy (anonymous):

7(d-5)+12=5 7d-35+12=5 7d-23=5 7d=28 d=4